OpenStudy (anonymous):

You have two lots to pick from. The budget for your lot is $160,000. Lot A requires a down-payment of $17,500 and cost $80 per square foot. The size of lot A is 30 feet by 70 feet. Lot B requires a down-payment of $20,000 and costs $70 per square foot. The size of lot B is 25 feet by 75 square feet. Based on your analysis of both lots, which one can you afford?

OpenStudy (anonymous):

The answer is B. If you need a demonstration tell me.

OpenStudy (anonymous):

A = 17,500 + 80 * 30 * 70 = 185,500 B = 20,000 + 70 * 25 * 75 = 151,250